the first term of an AP is 11 and the sums of the first five and first ten terms respectively are equal in magnitude but opposite in sign. find the second term of the progression.

Answer by mananth(16946)

the first term of an AP is 11 and the sums of the first five and first ten terms respectively are equal in magnitude but opposite in sign. find the second term of the progression.
Hera a =11
Sn = (n/2)(2a+(n-1)d)
S5 = 5/2(2*11+(5-1)*d)
=5/2(22+4d))

S10 = 10/2(2*11+9*d)
=5(22+9d)
S5= -S10
5/2(22+4d) = -5(22+10d)
22+4d=-2(22+10d)
22+4d=-44-20d
24d= -66
d= -66/24
d=-11/4
T2 = 11+d
T2 = 11-11/4
T2=33/4
11, 33/4,22/4.11/4,0,-11/4, -22/4,-33/4,-44/4,-55/4,